ENUMERATING A_{3}(2) BLUEPRINTS, AND AN APPLICATION

MICHAEL ABRAMSON AND CURTIS D. BENNETT

Abstract. We present an enumeration scheme for all blueprints of an A_{2}(p) building (with p prime). We then provide a computer based proof that the Ã_{n}(ℚ_{2}) buildings do not conform to a blueprint for n ≥ 3.